# Boosted second moment method in random regular graphs

### v2

@@ 1. Introduction
In the Appendix we provide sharper estimates for $\alpha^{\mathrm{FM}}_{d}$, see Theorem 11.1.
In [11, 12] Frieze and Łuczak used a second moment argument to obtain an asymptotically matching lower bound to prove that $\alpha^{\star}_{d}$ is actually asymptotically equal to (1).
@@
Note, however, that neither the Frieze–Łuczak, nor the Ding–Sly–Sun result is helpful when one is interested in $\alpha^{\star}_{d}$ for some specific degree, say, for $d=100$. What concrete lower bounds do we have at our disposal? Most results in this direction attained lower bounds by analyzing certain (local) processes/algorithms. [^3] Many results actually apply to the more general setting of graphs of (essentially) large girth. A notable tool is Wormald’s differential equation method. [^4] This method was introduced in [21]. See [19] for a concise treatment. The difficulty with this method is that it tends to be very challenging to carry out the required computer analysis for more sophisticated algorithms. Due to such difficulties, some approaches treat only the smallest cases $d=3$ or $d=3,4$. Currently, the best lower bounds are due to Csóka [6] for $d=3,4$; to Duckworth–Zito [10] for $d=5,6,7$; and to Wormald [21] for $d\geq 8$, apart from $d=16$ for which degree a better bound was achieved via star decompositions in [7].
@@ 1.1. Our bounds
We will see that verifying condition (3) essentially requires comparing two local maxima of $f_{d,\alpha}$. Therefore, it is straightforward to determine $\alpha^{\mathrm{SM}}_{d}$ with high precision using a computer for a specific degree $d$. It turns out that the lower bound $\alpha^{\mathrm{SM}}_{d}$ would already provide new results for $d\geq 20$, but we can actually boost it further with the following additional step. We will see that the second moment method proves the existence of independent sets with a certain spatial Markov property. This means that if we make “local improvements” to our independent set, then we can precisely calculate the increased density. In Section 5 we will see a simple procedure that gives the following augmented bound:
@@
As we mentioned, although our bounds do not come in the shape of a closed formula, it is relatively simple to obtain precise numerical values with a computer. Our SageMath code is able to quickly compute the bound for any given degree $d\leq 10^{9}$. We uploaded the code to the following GitHub repository.
@@ 1.2. Applications
The significance of the existence of Markovian independent sets of density $\alpha=\alpha^{\mathrm{SM}}_{d}$ goes beyond obtaining the lower bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{aug}}_{d}$. One can start from a Markovian independent set and try to create other objects, and hence proving the existence of more complex structures in random regular graphs. (For instance, suppose that we want to find large induced subgraphs with degree bound, say, $2$. Then it is not hard to design a local algorithm that produces such a subgraph from an independent set.)
@@
  • In Section 2 we provide a proof sketch for Theorem 1.1. Later we will rigorously prove a stronger version (Theorem 4.3) building on a result of Backhausz–Bordenave–Szegedy [1] about typical processes. This section is not required for the rest of the paper, but it may help the reader to gain a clearer understanding.
@@ 2. Counting independent sets in random regular graphs
Our aim here is to present a short, self-contained proof of our basic bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{SM}}_{d}$ (Theorem 1.1). The argument relies only on standard results from the literature and makes direct use of the second moment method.
@@
By the Paley–Zygmund inequality we get
@@
(9) $\mathbb{P}(Z_{N}>0)>\frac{1}{(Nd)^{K}}\text{ for some constant }K.$
@@
As we mentioned in the introduction, the independence ratio of $\mathbb{G}_{N,d}$ is tightly concentrated around its expectation. We will use the following precise statement that readily follows from [20, Theorem 2.19], which is a general Azuma-type concentration result. Let $\alpha(G)$ denote the independence ratio of $G$; then
(10) $\mathbb{P}\bigg(\big|\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})-\mathbb{E}\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\big|\geq t\bigg)\leq 2\exp\left(\frac{-Nt^{2}}{d}\right).$
@@
Since the event $Z_{N}>0$ is equivalent to $\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\alpha$, from (9) and (10) we obtain
$\mathbb{E}\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\alpha-t_{N}.$
@@
$\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\mathbb{E}\alpha(\mathbb{G}^{\textrm{conf}}_{N,d})-t_{N}\geq\alpha-2t_{N}.$
@@ 3. Typical processes
When studying labelings or colorings of random regular graphs, a useful tool is to study processes on the limiting graph, which is the $d$-regular infinite tree $T_{d}$. It is well known that $\mathbb{G}_{N,d}$ converges to $T_{d}$ locally (Benjamini–Schramm convergence) as $N\to\infty$. This is due to the fact that $\mathbb{G}_{N,d}$ is essentially large-girth with high probability: for every $\ell\in\mathbb{N}$ the number of cycles of length at most $\ell$ is $o(N)$.

### v3

@@ 1. Introduction
We mention here that we will work out sharper estimates for $\alpha^{\mathrm{FM}}_{d}$ in the Appendix; see Theorem 11.1.
In [11, 12] Frieze and Łuczak used a second moment argument to obtain a lower bound matching (1). In conclusion, $\alpha^{\star}_{d}$ is asymptotically equal to (1).
@@
Note, however, that neither the Frieze–Łuczak, nor the Ding–Sly–Sun result is helpful when one is interested in $\alpha^{\star}_{d}$ for some specific degree, say, for $d=100$. What concrete lower bounds do we have at our disposal? Most results in this direction attained lower bounds by analyzing certain local processes or algorithms. [^3] Many results actually apply to the more general setting of graphs of (essentially) large girth. A notable tool is Wormald’s differential equation method. [^4] This method was introduced in [21]. See [19] for a concise treatment. The difficulty with this method is that it tends to be very challenging to carry out the required computer analysis for more sophisticated algorithms. Due to such difficulties, some approaches treat only the smallest cases $d=3$ or $d=3,4$. Currently, the best lower bounds are due to Csóka [6] for $d=3,4$; to Duckworth–Zito [10] for $d=5,6,7$; and to Wormald [21] for $d\geq 8$, apart from $d=16$ for which degree a better bound was achieved via star decompositions in [7].
@@ 1.1. Our bounds
We will see that verifying condition (3) essentially requires comparing two local maxima of $f_{d,\alpha}$. Therefore, it is straightforward to determine $\alpha^{\mathrm{SM}}_{d}$ with high precision using a computer for a specific degree $d$. It turns out that the lower bound $\alpha^{\mathrm{SM}}_{d}$ would already provide new results for $d\geq 20$, but we can actually boost it further with an additional step.
This boosting is made possible by a strengthened version of Theorem 1.1, which guarantees the existence of an independent set (of density $\alpha^{\mathrm{SM}}_{d}$) satisfying a certain spatial Markov property (see Theorem 4.3). The argument relies on a refinement of the second moment method, using a result of Backhausz–Bordenave–Szegedy [1] on typical processes (see Lemma 3.3). Exploiting the Markov property, we perform local improvements to the independent set and compute explicitly the resulting gain in density. In Section 5 we present a simple procedure that yields the following augmented bound:
@@
As we mentioned—even though our bounds do not come in the shape of a closed formula—it is relatively simple to obtain precise numerical values with a computer. Our SageMath code is able to quickly compute the bound for any given degree $d\leq 10^{9}$. We uploaded the code to the following GitHub repository.
@@ 1.2. Applications
The significance of the existence of Markovian independent sets of density $\alpha=\alpha^{\mathrm{SM}}_{d}$ goes beyond obtaining the lower bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{aug}}_{d}$. One can start from a Markovian independent set and try to create other objects, and hence proving the existence of more complex structures in random regular graphs.
@@
  • In Section 2 we provide a proof sketch for Theorem 1.1. Later we will rigorously prove a stronger version (Theorem 4.3) building on a result of Backhausz–Bordenave–Szegedy [1] about typical processes. Although this section is not required for the rest of the paper, it may help the reader gain a clearer understanding, and it outlines a direct path to proving Theorem 1.1.
@@ 2. Counting independent sets in random regular graphs
Our aim here is to outline a short, self-contained proof of our basic bound $\alpha^{\star}_{d}\geq\alpha^{\mathrm{SM}}_{d}$ (Theorem 1.1). The argument relies only on standard results from the literature and makes direct use of the second moment method.
@@
By the Paley–Zygmund inequality [^8] In this case this is simply the Cauchy–Schwarz inequality for the product $\mathbbm{1}_{\{Z_{N}>0\}}Z_{N}$. we get
@@
(9) $\mathbb{P}(Z_{N}>0)\geq\frac{1}{(Nd)^{K}}\text{ for some constant }K.$
@@
As we mentioned in the introduction, the independence ratio of $\mathbb{G}_{N,d}$ is tightly concentrated around its expectation. We will use the following precise statement that readily follows from [20, Theorem 2.19], which is a general Azuma-type concentration result. Let $i(G)$ denote the independence ratio of $G$; then
(10) $\mathbb{P}\bigg(\big|i(\mathbb{G}^{\textrm{conf}}_{N,d})-\mathbb{E}\,i(\mathbb{G}^{\textrm{conf}}_{N,d})\big|\geq t\bigg)\leq 2\exp\left(\frac{-Nt^{2}}{d}\right).$
@@
Since the event $Z_{N}>0$ is equivalent to $i(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\alpha$, from (9) and (10) we obtain
$\mathbb{E}\,i(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\alpha-t_{N}.$
@@
$i(\mathbb{G}^{\textrm{conf}}_{N,d})\geq\mathbb{E}\,i(\mathbb{G}^{\textrm{conf}}_{N,d})-t_{N}\geq\alpha-2t_{N}.$
@@ 3. Typical processes
When studying labelings or colorings of random regular graphs, a useful tool is to study processes on the limiting graph, which is the $d$-regular infinite tree $T_{d}$. It is well known that $\mathbb{G}_{N,d}$, as well as $\mathbb{G}^{\textrm{conf}}_{N,d}$, converges to $T_{d}$ locally (Benjamini–Schramm convergence) as $N\to\infty$. This is due to the fact that $\mathbb{G}_{N,d}$ is essentially large-girth with high probability: for every $\ell\in\mathbb{N}$ the number of cycles of length at most $\ell$ is $o(N)$.